OpenStudy (anonymous):

I need help please:

OpenStudy (anonymous):

the questions is:

OpenStudy (anonymous):

2(e^2-e)

OpenStudy (anonymous):

can u please show me how u got the answer?

OpenStudy (anonymous):

make substitution sqrt(x)=y

OpenStudy (anonymous):

yes i reach to the point where i get (e^u)/u but then i dont know how to integrate this

OpenStudy (anonymous):

\[\sqrt{x}=y \implies dx=2dy.\] \[\int\limits_{1}^{2}2e^ydy=2(e^2-e)\]

OpenStudy (anonymous):

ooookay, i get it now.... Thanks :]

OpenStudy (anonymous):

sorry, it will be dx=2ydy

OpenStudy (anonymous):

well its good you understood;)